Experiment 11: Investigating Bandgap Energies, Materials, and Design of Light-Emitting Diodes (LEDs) 4. The band gap in aluminum phosphide (AIP) is 2.5 electronvolts. What wavelength of light is emitted by an AIP diode? Show work. 5. An aluminum antimonide solid-state laser emits light with a wavelength of 730 nm. Calculate the band gap in joules. Show work. 6. Explain the difference between the intrinsic band gap and the active or optical band gap.
Answer #1

Answer:

4) The relation between band gap (energy) and wavelength is

Wavelength(in nm)=1240.8/Energy(in eV)

Given band gap of Aluminum phosphide=2.5 eV,

therefore Wavelength(in nm)=1240.8/Energy(in eV)

= 1240.8/2.5=496.32 nm

Wavelength(in nm)=496.32 nm.

(5) Given Wavelength(in nm)=730 nm,

then Wavelength(in nm)=1240.8/Energy(in eV)

Energy(in eV) =1240.8/Wavelength(in nm)=1240.8/730 nm=1.699 eV

Since 1 eV=1.6x10-19 J, then 1.699 eV=1.699x1.6x10-19 J

Band gap Energy=2.72x10-19 J.

As per our guidelines I have to do only first question but I answered two for you. Please like this answer. Thank you.